# Bias-Field Digitized Counterdiabatic Quantum Algorithm for Higher-Order Binary Optimization

### v1

@@ Abstract
We present an enhanced bias-field digitized counterdiabatic quantum optimization (BF-DCQO) algorithm to address higher-order unconstrained binary optimization (HUBO) problems. Combinatorial optimization plays a crucial role in many industrial applications, yet classical computing often struggles with complex instances. By encoding these problems as Ising spin glasses and leveraging the advancements in quantum computing technologies, quantum optimization methods emerge as a promising alternative. We apply BF-DCQO with an enhanced bias term to a HUBO problem featuring three-local terms in the Ising spin-glass model. Our protocol is experimentally validated using 156 qubits on an IBM quantum processor with a heavy-hex architecture. In the studied instances, the results outperform standard methods, including the quantum approximate optimization algorithm (QAOA), quantum annealing, simulated annealing, and Tabu search. Furthermore, we perform an MPS simulation and provide numerical evidence of the feasibility of a similar HUBO problem on a 433-qubit Osprey-like quantum processor. Both studied cases, the experiment on 156 qubits and the simulation on 433 qubits, can be considered as the start of the commercial quantum advantage era, Kipu dixit, and even more when extended soon to denser industry-level HUBO problems.
I Introduction
@@
Combinatorial optimization problems arise in a multitude of situations in science and industry, from logistics and scheduling to computational chemistry and biology. In combinatorial optimization, the best or near-optimal solution is searched within a finite but large discrete configuration space. Theoretical tools from statistical physics have given insights into these problems, as their connection to disordered systems was identified early on Ref. [1]. These problems generally belong to the NP-hard complexity class, requiring a computation time that grows exponentially with the problem size on classical computers. While complex problem instances test the limits of classical computing, quantum computing has emerged as a suitable counterpart to drive the state-of-the-art in this field. This is brought forward by the rapid development of quantum technologies combined with the natural mapping of combinatorial optimization problems to the widely studied Ising spin glass model, whose ground state encodes the optimal solution [2]. Several methods have been proposed so far, including adiabatic quantum optimization [3] and the QAOA [4], which can be thought as a digitized version of an adiabatic quantum computing procedure [5].
@@
In addition to the capacity of quantum hardware to embed large-scale systems, the potential of quantum computing to surpass classical optimization methods has been demostrated through dedicated quantum algorithms [6, 7, 8, 9, 10, 11, 12, 13]. Among these, QAOA has emerged as one of the front-runner candidates for showing quantum speedup in optimization. Recent studies on 8-SAT up to 20 variables [14] show that the time to solution for QAOA outperforms best-known classical methods. Similar evidence of quantum speedup for solving the low autocorrelation binary sequence (LABS) problem has also been found [15, 16, 17].
@@
In parallel with quantum algorithms, quantum-inspired algorithms run on classical hardware, especially those based on tensor networks (TNs) [18, 19, 20, 21, 22], have evolved rapidly in the recent years. They were developed over the past three decades mostly for the study of condensed-matter physics [23, 21] and have been adopted as efficient simulators of quantum hardware, as classical benchmarks for quantum algorithms, and in hybrid approaches combining TNs and quantum algorithms [24, 25]. The recent kicked-Ising-model experiment of IBM on a 127-qubit device [26, 27, 28, 29, 30] and D-Wave’s simulation of quantum quench dynamics across a spin-glass phase transition [31, 32] highlighted the interplay of quantum simulations and TNs. For combinatorial optimization, new approaches have combined generative machine learning models with TNs [33, 34, 35], and ground-state search methods have been applied to problems with up to quadratic terms on higher-dimensional graphs [36, 37, 38].
@@
Compared to quadratic problems, higher-order ones are significantly more difficult to solve. While a number of experiments have recently demonstrated HUBO implementations on quantum hardware [39, 40, 41], there are hardware limitations that complicate the effective implementation of these protocols on large- and even intermediate-scale problems. These include short decoherence times, qubit connectivities, and the presence of noise. Several alternative methods, such as counterdiabatic driving (CD) protocols, have been proposed to overcome these issues [42, 43, 44, 45, 46, 47]. By suppressing diabatic transitions, CD protocols can find better solutions in different scenarios.
@@
In this work, we exploit the recently developed BF-DCQO formulation [48] to address HUBO problems on both ideal simulators and hardware. Here, we make use of the entire new 156-qubit IBM quantum computer ibm_fez [49] and a matrix product state (MPS) simulation of the upcoming IBM Osprey processor with 433 qubits. All results are compared with those obtained through well-established classical optimization methods as well as the D-Wave quantum annealing platform [50], known for solving quadratic unconstrained binary optimization (QUBO) problems naturally because of its structure. For both methods, extra qubits are needed for mapping HUBO problems into implementable QUBO instances. We furthermore obtain a reference MPS solution using the density matrix renormalization group (DMRG) method for ground-state search [51]. Even though our work addresses problems with up to three-body terms, our main findings can be extended to higher-order optimization problems, including the LABS problem [17], MAX $k$-SAT [52], factorization [53] and protein folding [54, 55], among others.
II Formulation
II.1 Digitized counterdiabatic quantum optimization (DCQO)
@@
Adiabatic quantum optimization aims to prepare the ground state of a given problem Hamiltonian. An initial state is evolved within the time window $t\in[0,T]$ under the adiabatic protocol $H_{\text{ad}}(\lambda)=(1-\lambda)H_{i}+\lambda H_{f}$. Here, $\lambda(t)$ is a time-dependent scheduling function that describes the adiabatic evolution of the system, starting from an initial Hamiltonian $H_{i}$, whose corresponding ground state is easy to prepare, towards the problem Hamiltonian $H_{f}$. Relying on the adiabatic theorem, this protocol reaches the desired $H_{f}$ ground state in the adiabatic limit $\dot{\lambda}(t)\to 0$. Following Ref. [48], we extend a disordered Ising Hamiltonian to up to three-body terms:
$H_{f}=\sum_{i}h_{i}^{z}\sigma^{z}_{i}+\sum_{ij}J_{ij}\sigma^{z}_{i}\sigma^{z}_{j}+\sum_{ijk}K_{ijk}\sigma^{z}_{i}\sigma^{z}_{j}\sigma^{z}_{k}.$ (1)
@@
The ground state of this Hamiltonian is called a $p$-spin glass with $p=3$ [56], and the model is commonly known in quantum optimization as the higher-order Ising chain [39, 40]. In the following, we use $\lambda(t)=\sin^{2}(\pi\sin^{2}(\pi t/2T)/2)$ as the scheduling function. We will use as initial Hamiltonian
@@
with $h_{j}^{x}$ ($h_{j}^{b}$) the transverse- (longitudinal bias-) field contributions acting on the $j$th spin. We set $h_{j}^{x}=-1$ and $h^{b}_{j}=0$, so that the $N$-qubit ground state of $H_{i}$ is
|$\ket{+}^{\otimes N}=\left(\frac{\ket{0}+\ket{1}}{\sqrt{2}}\right)^{\otimes N}.$ (3)
@@
where $\mathcal{O}_{0}(t)=\partial_{\lambda}H_{\text{ad}}$ and $\mathcal{O}_{k}(t)=[H_{\text{ad}},\mathcal{O}_{k-1}(t)]$. In the limit $l\to\infty$, this expansion converges to the exact gauge potential. The coefficients $\alpha_{k}$ are obtained by minimizing the action $S_{l}=\text{tr}[G_{l}^{2}]$ with $G_{l}=\partial_{\lambda}H_{\text{ad}}-i\big{[}H_{\text{ad}},A^{(l}_{\lambda}\big{]}$. For simplicity, we consider the first-order ($l=1$) nested-commutator term (see Appendix A for a step-by-step derivation), which takes the form
@@
In the fast evolution regime, the adiabatic term $H_{\text{ad}}(\lambda)$ in Eq. (4) can be omitted, reducing the number of required quantum gates and thereby decreasing hardware noise, while still maintaining the quality of the solution [62, 55, 48, 63]. This approach is adopted in the present study, where only the CD contribution is considered without compromising performance. The time evolution of such Hamiltonians remains a challenge on current analog quantum platforms, and may even be highly difficult due to the nonstoquasticity [64]. To address this issue, digitized counterdiabatic quantum protocols have been proposed for implementation on digital quantum computers [65]. The resulting time-evolution operator can be decomposed using a first-order Trotter-Suzuki product formula [66] and digitized under a gate-based model approach as follows
@@
where $H_{\text{cd}}(t)=\sum_{j=1}^{n_{\text{terms}}}\gamma_{j}(t)H_{j}$ is encoded in $n_{\text{terms}}$ different Pauli operators $H_{j}$ and $n_{\text{trot}}$ are the Trotter steps considered with $\Delta t=T/n_{\text{trot}}$. The number of gates required for implementation is reduced by optimally decomposing the desired circuit following Refs. [67, 68]. Furthermore, we set a gate-cutoff threshold $\theta_{\text{cutoff}}$ in our experiments, such that any term in absolute value of (7) with an angle magnitude below this threshold is discarded, thus if $|\gamma_{j}(k\Delta t)\Delta t|\text{ mod }2\pi<\theta_{\text{cutoff}}$ is met. The larger the $\theta_{\text{cutoff}}$, the lower is the amount of resources that are required, but the obtained circuit expressivity is also lower. Therefore, it is crucial to select an appropriate value of this parameter to significantly reduce the amount of gates to implement without compromising the quality of the outcomes.
II.2 Bias-field updating protocol
@@
In BF-DCQO, the standard DCQO is performed iteratively, where the solution from each iteration serves as the input bias for the next iteration [47, 48, 69]. At each step, the bias field $h^{b}_{i}$ for the $i$th qubit is updated based on the measured longitudinal component of the qubit, $\braket{\sigma^{z}_{i}}$. This update follows the rule $h^{b}_{i}=h^{b}_{i}(\braket{\sigma^{z}_{i}})$, where the functional form of $h^{b}_{i}$ represents the specific bias-field strategy employed. In this work, we explore the following strategies:
@@
  • Unsigned antibias($+$)/bias(--): $h^{b}_{i}=\pm\braket{\sigma^{z}_{i}}$. This strategy offers bias fields whose strength is exactly the longitudinal component. This is a sensible choice when a qubit is effectively close to an equal superposition, such that there is no clear direction of bias.
@@
  • Signed antibias($+$)/bias(--): $h^{b}_{i}=\pm\operatorname{sgn}\braket{\sigma^{z}_{i}}$. This strategy offers bias fields that are equivalent to the effective orientation of the qubit, regardless the magnitude. This is a sensible choice when a qubit has a clear preferred direction of bias.
@@
Once a suitable updating method has been selected, BF-DCQO iteratively measures the final state after time evolution in the computational basis and the longitudinal bias fields $h^{b}_{i}$ are updated accordingly. In the following, we set $h^{b}_{i}=0$ for the first bias-field iteration. According to the findings of Ref. [48], modifying the selected updating strategy throughout the entire routine can circumvent local minima and lead to faster convergence. For our results, we use the unsigned bias followed by a final iteration with a weighted signed bias. Specifically, the strength of $h^{b}_{i}$ is increased by rescaling $h^{b}_{i}\mapsto\text{constant}\times h^{b}_{i}$. In our experiments, we rescale by a factor of five to ensure that the magnitude of the bias is significantly larger than the transverse term, restricting the search space near the previous solution. Since the solution provided by DCQO, from the first iteration, samples low-energy states of the spin-glass Hamiltonian [Eq. (1)], the expectation values of local Pauli-$z$ operators can be used as an effective bias in subsequent iterations, facilitating convergence.
The initial Hamiltonian is changed after each bias-field update, so that its ground state also changes and has to be prepared for the circuit execution. The lowest eigenvalue of each $h_{i}^{x}\sigma^{x}_{i}-h^{b}_{i}\sigma^{z}_{i}$ term of $H_{i}$ is given by $\lambda^{\min}_{i}=-\sqrt{(h^{b}_{i})^{2}+(h^{x}_{i})^{2}}$ with eigenstate |$\ket{\tilde{\phi}}_{i}=[h^{x}_{i},h^{b}_{i}+\lambda^{\min}_{i}]^{\text{T}}$ up to normalization. This product state can be efficiently prepared in a depth-one circuit layer by means of single-qubit $y$-axis rotations such that
|$\ket{\tilde{\psi}}=\bigotimes_{i=1}^{N}\ket{\tilde{\phi}}_{i}=\bigotimes_{i=1}^{N}R_{y}(\theta_{i})\ket{0}_{i}$ (8)
with $\theta_{i}=\tan^{-1}(h^{x}_{i}/(h^{b}_{i}-\lambda^{\min}_{i}))$.
@@
Once the circuit has been prepared, all qubits are measured $n_{\text{shots}}$ times, returning an $N$-bitstring $x_{0}\dots x_{N-1}$. Each measurement trivially translates into a sample of $H_{f}$ due to its construction as a weighted sum of Pauli-$z$ operators. Denoting each of these samples as $E_{k}$, the sample mean energy is defined as $E\coloneqq(1/n_{\text{shots}})\sum_{k=1}^{n_{\text{shots}}}E_{k}$. It can be used as an estimator of the expectation value of $H_{f}$ after ground-state preparation.
@@
It has been demonstrated that incorporating Conditional Value-at-Risk (CVaR) techniques on quantum optimization routines [73] can lead to significant improvements. Building upon this knowledge, in this work, we propose a CVaR-inspired method for iteratively updating the bias fields. Accordingly, let $\alpha$ be the ratio between the number of measurements considered and the total number of shots $n_{\text{shots}}$. Rather than taking the entire distribution of samples into account ($\alpha=1$), we first sort them in terms of energy, thus $E_{k}\leq E_{k+1}$. Next, the bias is updated using a fraction $10<\alpha<$ 1 of the lowest-energy outcomes taken from
@@
thus taking only a part of the total distribution into account. From the results drawn from both noiseless simulators and hardware, faster convergence to better solutions is expected for values of $\alpha\sim 1\%$.
@@
In all our studies, the reference ground-state energies $E_{0}$ are obtained by applying the DMRG algorithm [51] implemented in the ITensor library [74] and sampling over random initial MPSs (see Appendix B).
III Results
III.1 HUBO instances considered
In this section, we present the HUBO Hamiltonians that will be employed for testing BF-DCQO. Based on the gate error rates and on the coupling map connectivities of the selected platforms, we work with a nearest-neighbour three-body spin glass. In order to tackle industry-related use cases, we furthermore consider the weighted MAX 3-SAT problem [52]. To the best of our knowledge, this is the first work to tackle a binary satisfiability problem with the aid of CD protocols. For both cases, the corresponding circuit decomposition is shown in Fig. 5 of Appendix C.
@@ III.1 HUBO instances considered
where $\braket{\cdot}$ indicates that the enclosed qubit indices are nearest neighbours (thus $j=i+1$ and $k=i+2$). The coupling constants $h_{i}$, $J_{ij}$ and $K_{ijk}$ are taken as random Gaussian-distributed numbers with zero mean and unit variance. The first-order nested-commutator expansion for this Hamiltonian is given by Eq. (6) but it only contains the nearest-neighbour qubit indices as in Eq. (12). Due to its linearity, it perfectly sticks to the heavy-hex IBM qubit coupling maps. Furthermore, the ease with which this and related models can be simulated using tensor networks makes their simulation an ideal benchmark for evaluating the performance of quantum platforms. Additionally, we conduct a similar test considering a 433-qubit NN HUBO instance, with the objective of simulating ideally (without noise) the expected performance of the upcoming IBM Osprey quantum platform via MPS. For this case, the values $n_{\text{shots}}=10000$, $\alpha=0.01$, $\theta_{\text{cutoff}}=0.06$ are set. The results can be seen in Fig. 1, with the amount of resources needed per iteration shown in Table 1.
Weighted MAX 3-SAT problem.—The maximum $k$-satisfiability (MAX $k$-SAT) problem belongs to the NP-complete complexity class [75] and its weighted variant is defined as follows: given a Boolean formula in conjunctive normal form (thus conjunction of disjuncted clauses), the weighted maximum $k$-satisfiability problem (MAX W-$k$-SAT) aims to find an assignment of truth-valued variables that maximizes the sum of weights of all the $k$-variable clauses met [52]. The MAX $k$-SAT problem can be recovered simply by setting the weights be equal to one. If we let $k=3$, this problem can be written as
@@
Here, $\omega_{ijk}$ is chosen as a random uniformly distributed variable between zero and one. As it is a maximization problem, it can be transformed into a minimization one by performing the map $H^{\text{MW3S}}_{f}\mapsto-H^{\text{MW3S}}_{f}$. The first-order nested-commutator of Eq. (15) yields

### v2

@@ Abstract
Combinatorial optimization plays a crucial role in many industrial applications. While classical computing often struggles with complex instances, quantum optimization emerges as a promising alternative. Here, we present an enhanced bias-field digitized counterdiabatic quantum optimization (BF-DCQO) algorithm to address higher-order unconstrained binary optimization (HUBO). We apply BF-DCQO to a HUBO problem featuring three-local terms in the Ising spin-glass model, validated experimentally using 156 qubits on an IBM quantum processor. In the studied instances, our results outperform standard methods such as the quantum approximate optimization algorithm, quantum annealing, simulated annealing, and Tabu search. Furthermore, we provide numerical evidence of the feasibility of a similar HUBO problem on a 433-qubit Osprey-like quantum processor. Finally, we solve denser instances of the MAX 3-SAT problem in an IonQ emulator. Our results show that BF-DCQO offers an effective path for solving large-scale HUBO problems on current and near-term quantum processors.
Introduction
@@
Combinatorial optimization problems arise in a multitude of situations in science and industry, from logistics and scheduling to computational chemistry and biology. In combinatorial optimization, the best or near-optimal solution is searched within a finite but large discrete configuration space. Theoretical tools from statistical physics have given insights into these problems, as their connection to disordered systems was identified early on [1]. These problems generally belong to the NP-hard complexity class, requiring a computation time that grows exponentially with the problem size on classical computers. While complex problem instances test the limits of classical computing, quantum computing has emerged as a suitable counterpart to drive the state of the art in this field. This is possible due to the rapid development of quantum technologies combined with the natural mapping of combinatorial optimization problems to the widely studied Ising spin glass model, whose ground state encodes the optimal solution [2]. Several methods have been proposed so far, including adiabatic quantum optimization [3] and the quantum approximate optimization algorithm (QAOA) [4], which can be thought as a digitized version of adiabatic quantum computing [5].
@@
In addition to the capacity of quantum hardware to embed large-scale systems, the potential of quantum computing to surpass classical optimization methods has been demostrated through dedicated quantum algorithms [6, 7, 8, 9, 10, 11, 12, 13]. Among these, QAOA has emerged as one of the front-runner candidates for showing quantum speedup in optimization. Recent studies on 8-SAT problems with up to 20 variables [14] show that the time to solution for QAOA outperforms best-known classical methods. Similar evidence of quantum speedup for solving the low autocorrelation binary sequence (LABS) problem has also been found [15, 16, 17].
@@
In parallel with quantum algorithms, quantum-inspired algorithms run on classical hardware, especially those based on tensor networks (TNs) [18, 19, 20, 21, 22], have evolved rapidly in the recent years. They were developed over the past three decades mostly for the study of condensed-matter physics [23, 21] and have been adopted as efficient simulators of quantum hardware, as classical benchmarks for quantum algorithms, and in hybrid approaches combining TNs and quantum algorithms [24, 25, 26, 27]. The recent kicked-Ising-model experiment of IBM on a 127-qubit device [28, 29, 30, 31, 32] and D-Wave’s simulation of quantum quench dynamics across a spin-glass phase transition [33, 34] highlighted the interplay of quantum simulations and TNs. For combinatorial optimization, new approaches have combined generative machine learning models with TNs [35, 36, 37], and ground-state search methods have been applied to problems with up to quadratic terms on higher-dimensional graphs [38, 39, 40].
@@
Compared to quadratic problems, higher-order ones are significantly more difficult to solve. While a number of experiments have recently demonstrated higher-order unconstrained binary optimization (HUBO) implementations on quantum hardware [41, 42, 43, 44], there are hardware limitations that complicate the effective implementation of these protocols on large- and even intermediate-scale problems. These include short decoherence times, qubit connectivities, and the presence of noise. Several alternative methods, such as counterdiabatic driving (CD) protocols, have been proposed to overcome these issues [45, 46, 47, 48, 49, 50]. By suppressing diabatic transitions, CD protocols can find better solutions in different scenarios.
@@
In this work, we develop and exploit an improved version of the recently introduced bias-field digitized counterdiabatic quantum optimization (BF-DCQO) formulation [51] to address HUBO problems on both ideal simulators and hardware. We make use of the entire new 156-qubit IBM quantum computer ibm_fez and a matrix product state (MPS) simulation of the upcoming IBM Osprey processor with 433 qubits. To test BF-DCQO on denser instances, we consider a MAX 3-SAT problem at the critical density and employ the 25-qubit IonQ Aria 1 emulator with a noise model. All results are compared with those obtained through well-established classical optimization methods as well as the D-Wave quantum annealing platform. Unlike these methods, BF-DCQO does not require additional qubits for mapping the initial HUBO problem into a quadratic one. Even though our work addresses problems with up to three-body terms, our method is equally applicable to higher-order optimization problems, including the LABS problem [17], MAX $k$-SAT [52], factorization [53] and protein folding [54, 55], among others.
Methods
Digitized counterdiabatic quantum optimization (DCQO)
Figure 1: Schematic of the BF-DCQO protocol. Given a binary optimization problem of arbitrary order (depicted in a) encoded as a $p$-spin glass [Eq. (1)], the BF-DCQO algorithm performs DCQO iteratively, updating the input bias $h_{j}^{b}$ from the solutions obtained from the previous step. Each iteration consists of running the corresponding DCQO circuit (depicted in b), sampling the final state, and updating the bias fields. It amounts to the use of (i) the CVaR-based approach, which considers only a subset of the low-energy samples, and (ii) the updating strategy as a function of the expectation value, i.e., $h^{b}_{j}(\braket{\sigma^{z}_{j}})$ (depicted in c).
@@ Methods
Adiabatic quantum optimization aims to prepare the ground state of a given problem Hamiltonian. An initial state is evolved within the time window $t\in[0,T]$ under the adiabatic protocol $H_{\text{ad}}(\lambda)=(1-\lambda)H_{i}+\lambda H_{f}$. Here, $\lambda(t)$ is a time-dependent scheduling function that describes the adiabatic evolution of the system, starting from an initial Hamiltonian $H_{i}$, whose corresponding ground state is easy to prepare, towards the problem Hamiltonian $H_{f}$. Relying on the adiabatic theorem, this protocol reaches the desired ground state of $H_{f}$ in the adiabatic limit $\dot{\lambda}(t)\to 0$. We consider the disordered Ising Hamiltonian with up to three-body terms:
$H_{f}=\sum_{i}h_{i}^{z}\sigma^{z}_{i}+\sum_{i<j}J_{ij}\sigma^{z}_{i}\sigma^{z}_{j}+\sum_{i<j<k}K_{ijk}\sigma^{z}_{i}\sigma^{z}_{j}\sigma^{z}_{k},$ (1)
@@
in contrast to Ref. [51], which only considers up to quadratic terms. We use here natural units with $\hbar=1$. The ground state of this Hamiltonian is called a $p$-spin glass with $p=3$ [56], and the model is commonly known in quantum optimization as the higher-order Ising chain [41, 42]. In the following, we use $\lambda(t)=\sin^{2}(\pi\sin^{2}(\pi t/2T)/2)$ as the scheduling function. The initial Hamiltonian is
@@
with $h_{j}^{x}$ ($h_{j}^{b}$) the transverse- (longitudinal bias-) field contributions acting on the $j$th spin. We set $h_{j}^{x}=-1$ and $h^{b}_{j}=0$, thus $H_{i}=-\sum_{j=0}^{N-1}\sigma^{x}_{j}$, so that the $N$-qubit ground state of $H_{i}$ is
[…]
@@
where $\mathcal{O}_{0}(t)=\partial_{\lambda}H_{\text{ad}}$ and $\mathcal{O}_{k}(t)=[H_{\text{ad}},\mathcal{O}_{k-1}(t)]$. In the limit $l\to\infty$, this expansion converges to the exact gauge potential. The coefficients $\alpha_{k}$ are obtained by minimizing the action $S_{l}=\text{tr}[G_{l}^{2}]$ with $G_{l}=\partial_{\lambda}H_{\text{ad}}-i\big{[}H_{\text{ad}},A^{(l}_{\lambda}\big{]}$. For simplicity, we consider the first-order ($l=1$) nested-commutator term (see Supplementary Note 1 for a step-by-step derivation), which takes the form
@@
In the fast evolution regime, the adiabatic term $H_{\text{ad}}(\lambda)$ in Eq. (4) can be omitted, reducing the number of required quantum gates and thereby decreasing hardware noise, while still maintaining the quality of the solution [62, 55, 51, 63]. This approach is adopted in the present study, where only the CD contribution is considered without compromising performance. The time evolution of such Hamiltonians remains a challenge on current analog quantum platforms, and may even be highly difficult due to the nonstoquasticity [64]. To address this issue, digitized counterdiabatic quantum protocols have been proposed for implementation on digital quantum computers [65]. The resulting time-evolution operator can be decomposed using a first-order Trotter-Suzuki decomposition [66] and digitized with a gate-based approach,
@@
where $H_{\text{cd}}(t)=\sum_{j=1}^{n_{\text{terms}}}\gamma_{j}(t)H_{j}$ is encoded in $n_{\text{terms}}$ different Pauli operators $H_{j}$ and $n_{\text{trot}}$ are the Trotter steps considered with $\Delta t=T/n_{\text{trot}}$. We use $T=1$ in units of the inverse energy scale relevant for each problem as defined below, and $n_{\text{trot}}=3$, thus one effective Trotter step is implemented. The number of gates required for implementation is reduced by optimally decomposing the desired circuit following Refs. [67, 68]. Furthermore, we set a gate-cutoff threshold $\theta_{\text{cutoff}}$ in our experiments, such that any term of Eq. (7) with an angle with absolute value below this threshold is discarded, thus if $|\gamma_{j}(k\Delta t)\Delta t|\text{ mod }2\pi<\theta_{\text{cutoff}}$ is met. The larger the $\theta_{\text{cutoff}}$, the lower is the amount of resources that are required, but the obtained circuit expressivity is also lower. Therefore, it is crucial to select an appropriate value of this parameter to significantly reduce the amount of gates to implement without compromising the quality of the outcomes.
Bias-field updating protocol
@@
In BF-DCQO, the standard DCQO is performed iteratively, so that the solution from each step serves as the input bias for the next iteration [50, 51, 69]. See Fig. 1 for a schematic diagram of the protocol. The initial Hamiltonian is changed after each bias-field update,
$\tilde{H}_{i}=H_{i}+\sum_{j=0}^{N-1}h^{b}_{j}(\braket{\sigma^{z}_{j}})\sigma^{z}_{j},$ (8)
[…]
@@
In each iteration, the bias field $h^{b}_{j}$ for the $j$th qubit is updated based on the measured longitudinal component of the qubit $\braket{\sigma^{z}_{j}}$. This update follows the rule $h^{b}_{j}=h^{b}_{j}(\braket{\sigma^{z}_{j}})$, where the functional form of $h^{b}_{j}$ represents the specific bias-field strategy employed. In this work, we explore the following strategies:
@@
  • Unsigned antibias(++)/bias(-): $h^{b}_{j}=\pm\braket{\sigma^{z}_{j}}$. This strategy offers bias fields whose strength is exactly the longitudinal component. This is a sensible choice when a qubit is effectively close to an equal superposition, such that there is no clear direction of bias.
@@
  • Signed antibias(++)/bias(-): $h^{b}_{j}=\pm\operatorname{sgn}\braket{\sigma^{z}_{j}}$. This strategy offers bias fields that are equivalent to the effective orientation of the qubit, regardless the magnitude. This is a sensible choice when a qubit has a clear preferred direction of bias.
@@
Once a suitable updating method has been selected, BF-DCQO iteratively measures the final state after time evolution in the computational basis and the longitudinal bias fields $h^{b}_{j}$ are updated accordingly. In the following, we set $h^{b}_{j}=0$ for the first bias-field iteration. According to the findings of Ref. [51], modifying the selected updating strategy throughout the entire routine can circumvent local minima and lead to faster convergence. For our results, we use the unsigned bias followed by a final iteration with a weighted signed bias. Specifically, the strength of $h^{b}_{j}$ is increased by rescaling $h^{b}_{j}\mapsto\kappa h^{b}_{j}$ with $\kappa>0$. In our experiments, we rescale by a factor of five to ensure that the magnitude of the bias is significantly larger than the transverse term, restricting the search space near the previous solution. Since the solution provided by DCQO, from the first iteration, samples low-energy states of the spin-glass Hamiltonian [Eq. (1)], the expectation values of local Pauli-$z$ operators can be used as an effective bias in subsequent iterations, facilitating convergence.
Although this updating strategy is employed in our experiments, alternative approaches might be more convenient depending on the nature of the problem at hand. Using $\kappa\in(0,1)$ grants more importance to the transverse contribution in $\tilde{H}_{i}$, whose ground state is an equal superposition of the computational basis. Thereby, a broader energy landscape is explored by flattening the expected sampling distribution. The algorithm is less prone to getting stuck in local minima with this choice, though it may significantly slow down convergence. On the other hand, using $\kappa>1$ provides greater strength to the longitudinal bias-field update, i.e. the optimal solution obtained from the last BF-DCQO iteration, constraining the search space around this solution. This accelerates the convergence but may lead to an undesired local minimum.
@@
Once the circuit has been executed, all qubits are measured $n_{\text{shots}}$ times, returning an $N$-bitstring $x_{0}\dots x_{N-1}$. Each measurement trivially translates into a sample of $H_{f}$ due to its construction as a weighted sum of Pauli-$z$ operators. Denoting each of these samples as $E_{k}$, the sample mean energy is defined as $E=(1/n_{\text{shots}})\sum_{k=1}^{n_{\text{shots}}}E_{k}$. It can be used as an estimator of the expectation value of $H_{f}$ after ground-state preparation.
@@
It has been demonstrated that incorporating Conditional Value-at-Risk (CVaR) techniques into quantum optimization routines [73] can lead to significant improvements. Recent studies have explored the impact of noise when sampling bitstrings from quantum computers. They determined upper and lower bounds for noiseless expectation values using CVaR noisy samples and quantified the sampling overhead needed to achieve good solutions [43]. Building upon this knowledge, we propose a CVaR-inspired method for iteratively updating the bias fields. Accordingly, let $\alpha$ be the ratio between the number of measurements considered and the total number of shots $n_{\text{shots}}$. Rather than taking the entire distribution of samples into account ($\alpha=1$), we first sort them in terms of energy so that $E_{k}\leq E_{k+1}$. Next, the bias is updated using a fraction $0<\alpha<1$ of the lowest-energy outcomes taken from
@@
thus taking only a part of the total distribution into account. From the results drawn from both noiseless simulators and hardware, we find the fastest convergence to the best solutions for values of $\alpha\sim 1\%$. This value is found empirically and the optimal value will depend on the problem and the number of shots. For example, we find that using smaller (larger) values of $\alpha$ shows faster convergence if the number of shots is large (small) with respect to the numbers used here.
@@
In all our studies of 156 and 433 qubits, the exact ground-state energies $E_{0}$ are obtained numerically by an iterative method similar to Ref. [74] (see Supplementary Note 2 for details). For 25-qubit instances, we obtain the exact solution using a brute force method.
HUBO instances considered
In this section, we present the HUBO Hamiltonians that are employed for testing BF-DCQO. Based on the gate error rates and on the coupling map connectivities of the selected platforms, we work with a nearest-neighbour three-body spin glass. In order to tackle industry-related use cases, we furthermore consider the weighted MAX 3-SAT problem [52]. To the best of our knowledge, this is the first work to tackle a binary satisfiability problem with the aid of CD protocols. For both cases, the corresponding circuit decomposition is shown in Supplementary Note 3.
@@
[…]
Weighted MAX 3-SAT problem.—The maximum $k$-satisfiability (MAX $k$-SAT) problem belongs to the NP-complete complexity class [76] and its weighted variant is defined as follows: given a Boolean formula in conjunctive normal form (thus conjunction of disjuncted clauses), the weighted maximum $k$-satisfiability problem (MAX W-$k$-SAT) aims to find an assignment of truth-valued variables that maximizes the sum of weights of all the $k$-variable clauses met [52]. The MAX $k$-SAT problem can be recovered simply by setting the weights equal to one. If we let $k=3$, this problem can be written as
@@
Here, $\omega_{ijk}$ is chosen as a random uniformly distributed variable between zero and one. As this is a maximization problem, it can be transformed into a minimization one by performing the map $H^{\text{MW3S}}_{f}\mapsto-H^{\text{MW3S}}_{f}$. The first-order nested-commutator of Eq. (16) yields